OpenStudy (kayders1997):

Can someone please help with integration by parts find the integral of x^2sinpixdx

OpenStudy (kayders1997):

@satellite73

OpenStudy (anonymous):

ack you have to do it twice

OpenStudy (anonymous):

\[u=x^2\\ du=2xdx\\ dv=\sin(\pi x)\\ v=-\frac{1}{\pi}\cos(\pi x)\] is hte first one

OpenStudy (kayders1997):

Omg

OpenStudy (kayders1997):

I don't like this already

OpenStudy (anonymous):

then you are going to have to do it again

OpenStudy (kayders1997):

Okay

OpenStudy (anonymous):

\[uv-\int v du\] is a start

OpenStudy (kayders1997):

Someone showed me a different method but we haven't learned it yet so it probably wouldn't be excepted on quiz/test

OpenStudy (kayders1997):

Okay I'll try and figure it out and I'll let you know what I get in a couple minutes

OpenStudy (anonymous):

\[-\frac{x^2}{\pi}\cos(\pi x)+\frac{2}{\pi}\int x\cos (\pi x)dx\]

OpenStudy (kayders1997):

Okay that's what I got

OpenStudy (daniel.ohearn1):

You can also do this one easy by the table method. Have you learned it yet?

OpenStudy (anonymous):

no

OpenStudy (daniel.ohearn1):

kayders?

OpenStudy (kayders1997):

No, I don't think that's what the other person showed me, I understood it way better though than this

OpenStudy (anonymous):

i mean no i have not learned it, but saw it in a movie with olmos

OpenStudy (kayders1997):

I meant I think

OpenStudy (kayders1997):

I don't like this method though to many things going on :(

OpenStudy (daniel.ohearn1):



OpenStudy (kayders1997):

Yes this was it

OpenStudy (daniel.ohearn1):

What I did was I chose x^2 to derive because eventually it becomes zero after deriving enough.. The other part is chosen to be the part I integrate.

OpenStudy (kayders1997):

Okay

OpenStudy (daniel.ohearn1):

now multiplying across the first line we have our original integral.. You'll notice that if we multiply diagonally though, each line down until we have 2 times the last integral row number we will have our answer taking into account the change in signs every row.

OpenStudy (kayders1997):

Oh okay

OpenStudy (kayders1997):

I like that method much more

OpenStudy (daniel.ohearn1):

It's a much more organized approach overall, when it works..

OpenStudy (kayders1997):

When it works?

OpenStudy (daniel.ohearn1):

When you don't need trig identities or other functional methods... Take the integral of 1/x(sqrt(x+1))dx

OpenStudy (kayders1997):

Oh my gosh :O

OpenStudy (kayders1997):

What calc are you in?

OpenStudy (daniel.ohearn1):

Integral calc

OpenStudy (kayders1997):

I'm a high school student in calc 1

OpenStudy (kayders1997):

There are many things I still need to learn lol

OpenStudy (kayders1997):

I'll try the problem you posted